A chef is using a mixture of two brands of italian dressing. the first brand contains 9% vinegar, and the second brand contains 14% vinegar. The chef wants to make 230 milimeters of a dressing that is 13% vinegar. How much of each brand should she use?
Answer by checkley77(12844)

.09x+.14(230-x)=.13*230
.09x+32.2-.14x=29.9
-.05x=29.9-32.2
-.05x=-2.3
x=-2.3/-.05
x=46 ml of 9% dressing.
230-46=184 ml of 14% dressing.
proof:
.09*46+.14*184=29.9
4.14+25.76=29.9
29.9=29/9